# An Intention-driven Lane Change Framework Considering Heterogeneous Dynamic Cooperation in Mixed-traffic Environment

### v1

@@ 1 Introduction
Modern lane change algorithms have evolved from classical models (Gipps (1986)) to contemporary data-driven approaches (Zheng (2014); Wang et al. (2019b)), and can be broadly categorized into three paradigms: rule-based, data-based, and incentive-based systems. Recent advances in inverse reinforcement learning (Sun et al. (2022)) and probabilistic modeling (Sun et al. (2018)) have partially incorporated driver intentions into decision-making (Sheng et al. (2022)). Nevertheless, several critical limitations persist. Firstly, existing frameworks often rely on a behavioral homogeneity assumption, which oversimplifies surrounding human drivers by adopting uniform behavioral models and thereby overlooks inter-driver heterogeneity in decision-making. Secondly, although some methods incorporate personalized intentions, they typically fail to capture the temporal evolution of interactions among surrounding vehicles. Lastly, there remains an intention modeling gap: current prediction approaches lack mechanisms to effectively link observable driving styles with latent reward structures, resulting in a disconnect between behavior recognition and motion-planning.
@@
To address the identified shortcomings in existing algorithms, we introduce an intention-driven lane change framework that synergistically integrates driving style recognition with adaptive decision-making as shown in Fig.1. In the context of this research, “driving style” is conceptualized as a habitual manner of vehicle operation, distinct to an individual or subset of drivers, that serves as a proxy for underlying personalized driving intentions. Our proposed lane change framework incorporates driving style into two distinct yet interrelated components: intention-prediction and motion-planning. The main contributions of this study are summarized as follows:
@@ 2.1 Automated Lane Change Algorithms
In summary, while each paradigm has advanced the modeling of lane change behavior, none fully resolves the challenge of safe and efficient AV–HV interactions in heterogeneous traffic. This motivates research that bridges interpretability, adaptability, and behavioral realism in unified frameworks.
@@ 2.3 Human Personality Integrated Lane Change Frameworks
To address this gap, we propose an intention-driven lane change framework that explicitly models the cooperation willingness of surrounding drivers in the target lane. By embedding both intrinsic and interactive personality-informed intentions into decision-making, the framework aims to achieve safer and more socially compliant AV maneuvers.
@@ 3 Problem Statement
Figure 2: Illustration of the NGSIM I-80 dataset: (a) Road segment layout; (b) Roadside camera view.
@@ 3.2 Dataset
We adopt the NGSIM I-80 dataset (Alexiadis et al. (2004)), illustrated in Fig. 2, for both training and evaluation. Collected via roadside cameras at a sampling rate of 10 Hz, the NGSIM dataset is widely regarded as one of the most comprehensive open-access repositories of naturalistic driving trajectories. It provides detailed vehicle-level information including position, speed, and lane assignment.
@@ 3.3 Lane Change Start-End Point Detection
Additionally, we sets forth a criterion to verify the continuous presence of lane lines, thereby reducing the likelihood of false detections. The efficacy of our detection algorithm is illustrated in Fig. 4, which highlights its capability to not only pinpoint rapid lane changes but also detect subtle lateral adjustments made by vehicles.

### v2

@@ 1 Introduction
Modern lane change algorithms have evolved from classical models (Gipps (1986)) to contemporary data-driven approaches (Zheng (2014); Wang et al. (2019b)), and can be broadly categorized into three paradigms: rule-based, data-driven, and incentive-based systems. Recent advances in inverse reinforcement learning (Sun et al. (2022)) and probabilistic modeling (Sun et al. (2018)) have partially incorporated driver intentions into decision-making (Sheng et al. (2022)). Nevertheless, several critical limitations persist. Firstly, existing frameworks often rely on a behavioral homogeneity assumption, which oversimplifies surrounding human drivers by adopting uniform behavioral models and thereby overlooks inter-driver heterogeneity in decision-making. Secondly, although some methods incorporate personalized intentions, they typically fail to capture the temporal evolution of interactions among surrounding vehicles. Lastly, there remains an intention modeling gap: current prediction approaches lack mechanisms to effectively link observable driving styles with latent reward structures, resulting in a disconnect between behavior recognition and motion-planning.
@@
To address the identified shortcomings in existing algorithms, we introduce an intention-driven lane change framework that synergistically integrates driving style recognition with adaptive decision-making as shown in Fig.1. In the context of this research, “driving style” is conceptualized as a habitual manner of vehicle operation, distinctive of an individual or subset of drivers, that serves as a proxy for underlying personalized driving intentions. Our proposed lane change framework incorporates driving style into two distinct yet interrelated components: intention-prediction and motion-planning. The main contributions of this study are summarized as follows:
@@ 2.1 Automated Lane Change Algorithms
In summary, rule-based methods offer interpretability but are limited by rigid assumptions, data-driven approaches achieve higher predictive accuracy but often lack interpretability and robustness, while incentive-based methods capture sequential decision-making but face challenges in reward design and scalability. Although both academic and industrial systems have demonstrated safe and efficient lane change performance in constrained domains, existing frameworks still struggle to accommodate heterogeneous and dynamically evolving human behaviors. Motivated by these limitations, we leverage the complementary strengths of data-based learning (for accurate pattern recognition) and incentive-based reasoning (for human-like reward modeling) to design an intention-driven lane change framework that explicitly incorporates human drivers’ cooperation intentions into the decision-making process.
@@ 2.3 Human Personality Integrated Lane Change Frameworks
To address this gap, we propose an intention-driven lane change framework that explicitly models the dynamic cooperation willingness of surrounding drivers in the target lane. By embedding both intrinsic and interactive personality-informed intentions into decision-making, the framework aims to achieve safer and more socially compliant AV maneuvers.
@@ 3.2 Dataset
This study employs the Next Generation SIMulation (NGSIM) I-80 dataset (Halkia and Colyar (2006)) for both training and evaluation. The dataset was collected on a section of the Interstate 80 freeway in California using multiple synchronized video cameras installed on nearby high-rise buildings. The recorded videos were subsequently processed to extract vehicle trajectories at a sampling rate of 10 Hz. The dataset provides detailed microscopic information such as vehicle position, velocity, and lane assignment, offering a naturalistic and high-resolution representation of mixed-traffic conditions. Owing to its comprehensiveness and accuracy, the NGSIM I-80 dataset has become one of the most widely adopted benchmarks for studying lane change behaviors and human–vehicle interactions.
@@ 3.3 Lane Change Start-End Point Detection
Additionally, we set forth a criterion to verify the continuous presence of lane lines, thereby reducing the likelihood of false detections. The efficacy of our detection algorithm is illustrated in Fig. 3, which highlights its capability to not only pinpoint rapid lane changes but also detect subtle lateral adjustments made by vehicles.